# Scaling Open-Source Multimodal Data by 10x: A Multimodal Dataset with One Trillion Tokens

### v1

@@ 5.1 Multimodal Interleaved Data
Large-scale multimodal interleaved datasets were first presented in Flamingo [Alayrac et al., 2022] and CM3 [Aghajanyan et al., 2022]. Kosmos [Huang et al., 2023] showed similar properties and was followed by Multimodal-C4 [Zhu et al., 2023] and OBELICS [Laurençon et al., 2023], the first open-source multimodal interleaved datasets. More recent work such as Chameleon [Meta, 2024] and MM1 [McKinzie et al., 2024] have scaled OBELICS to train state-of-the-art multimodal models. A complementary line of work, Mantis [Jiang et al., 2024] and MIMIC-IT [Li et al., 2023] builds interleaved instruction tuning datasets.
@@ 5.2 Large Open-source Pre-training Datasets
Large, high-quality pre-training datasets are the backbone of open-source research. In image-text datasets, where preliminary works [Schuhmann et al., 2021, Byeon et al., 2022, Schuhmann et al., 2022, Gadre et al., 2023] focused on scaling image-text datasets to billions of samples and has been crucial for training strong open-source multimodal models [Ilharco et al., 2021, Sun et al., 2023a]. Similarly in language modeling, large datasets like Pile [Gao et al., 2020], Redpajama [Computer, 2023], RefinedWeb [Penedo et al., 2023], Dolma [Soldaini et al., 2024], and FineWeb [Penedo et al., 2024] have been crucial for training fully transparent open-source models.
@@ 6 Limitations and conclusion
In this work, we present MINT-1T, the first open-source trillion token multimodal interleaved dataset and an important component for training large multimodal models. We believe MINT-1T will be a valuable resource for the research community to do open science on multimodal interleaved datasets. An important consideration when releasing large datasets is to avoid exposing harmful content from the web. We were careful about filtering out personally identifiable information and not safe for work content from MINT-1T. However, future work should explore more ways to improve the safety of multimodal internet data. Moreover, subsequent work should train models on larger subsets of MINT-1T, build novel multimodal document filtering methods to improve data quality, and curate multimodal sequences from other sources.
@@
We thank Jeffrey Li and Alex Fang for helpful discussions regarding data filtering and deduplication, Irena Gao for leading the development of the new OpenFlamingo training codebase which we use, Honglu Zhou for maintaining the evaluation code, and Srinath Meadusani and Lavanya Karanam for infrastructure support.
@@ A.2 Evaluation Details
We ablated multiple prompts for captioning and visual question answering benchmarks and found prompts that worked best across all Phi-3 mini based large multimodal models. For COCO captioning we used the prompt "A short description of this image in one sentence:". As TextCaps is a more OCR intensive captioning task we used "A short description of this image that describes visual content and explicitly spells out any text it contains:". For VQAv2, VizWiz, and OKVQA we used the prompt "Question: <question> Short answer in one phrase or single word:" and for TextVQA we used "Question: <question> Short answer in a very single phrase:". For all tasks we used greedy decoding with a maximum generation length of 56 tokens.
@@ A.3.1 Motivation
    The dataset is created by a team from the University of Washington, Salesforce Research, Stanford University, and University of California Berkeley.
@@ A.3.2 Composition
    In total there are 927.6 million (M) documents (913M HTML documents, 14M PDF documents, and 0.6M ArXiv documents).
@@
    MINT-1T’s HTML documents are filtered from CommonCrawl WARC dumps from 2013 to 2023 based on text quality, the presence of duplicated and undesirable content, and availability of images for downloading. MINT-1T PDF documents are filtered from CommonCrawl WAT dumps from 2023 to 2024 also based on text quality and the presence of duplicated and undesirable content. MINT-1T ArXiv documents is a subset of all ArXiv documents.
@@
    MINT-1T is not self-contained and relies on downloading external image and PDF urls to collect the full dataset. There are no guarantees these urls will remain available over time as link rot is a common problem for large scale image datasets. Moreover, as this dataset is over 300TB large, it is infeasible for us to host the full datasets (with images included). There are no restrictions regarding downloading images/PDFs from these external urls.

### v2

@@ 5.1 Multimodal Interleaved Data
Large-scale multimodal interleaved datasets were first presented in Flamingo [Alayrac et al., 2022] and CM3 [Aghajanyan et al., 2022]. Kosmos [Huang et al., 2023] showed similar properties and was followed by Multimodal-C4 [Zhu et al., 2023] and OBELICS [Laurençon et al., 2023], the first open-source multimodal interleaved datasets. More recent work such as Chameleon [Meta, 2024] and MM1 [McKinzie et al., 2024] have scaled OBELICS to train state-of-the-art multimodal models. A complementary line of work, Mantis [Jiang et al., 2024] and MIMIC-IT [Li et al., 2023] builds interleaved instruction tuning datasets. Similarly, Multimodal Arxiv [Li et al., 2024a] builds high quality captioning and instruction tuning data from ArXiv papers.
@@ 5.2 Large Open-source Pre-training Datasets
Large, high-quality pre-training datasets are the backbone of open-source research. In image-text datasets, where preliminary works [Schuhmann et al., 2021, Byeon et al., 2022, Schuhmann et al., 2022, Gadre et al., 2023] focused on scaling image-text datasets to billions of samples and has been crucial for training strong open-source multimodal models [Ilharco et al., 2021, Sun et al., 2023a]. Similarly in language modeling, large datasets like Pile [Gao et al., 2020], Redpajama [Computer, 2023], RefinedWeb [Penedo et al., 2023], Dolma [Soldaini et al., 2024], Datacomp-LM [Li et al., 2024b], and FineWeb [Penedo et al., 2024] have been crucial for training fully transparent open-source models.
@@ 6 Limitations and conclusion
In this work, we present MINT-1T, the first open-source trillion token multimodal interleaved dataset and an important component for training large multimodal models. We believe MINT-1T will be a valuable resource for the research community to do open science on multimodal interleaved datasets. An important consideration when releasing large datasets is to avoid exposing harmful content from the web. We were careful about filtering out personally identifiable information and not safe for work content from MINT-1T. However, future work should explore more ways to improve the safety of multimodal internet data. Moreover, subsequent work should train models on larger subsets of MINT-1T, build filtering methods to improve data quality, and curate multimodal sequences from other sources.
@@
We thank Srinath Meadusani and Lavanya Karanam for working on infrastructure, Jeffrey Li and Alex Fang for helpful discussions regarding data filtering and deduplication, Irena Gao for leading the development of the new OpenFlamingo training codebase which we use, Honglu Zhou for maintaining the evaluation code, James Park and Marianna Nezhurina for helpful discussions regarding PDF parsing, and Paul Josel for helping us with figure design.
@@ A.2 Evaluation Details
We ablated multiple prompts for captioning and visual question answering benchmarks and found prompts that worked best across all Phi-3 mini based large multimodal models. For COCO captioning we used the prompt "A short description of this image in one sentence:". As TextCaps is a more OCR intensive captioning task we used "A short description of this image that describes visual content and explicitly spells out any text it contains:". For VQAv2, VizWiz, and OKVQA we used the prompt "Question: <question> Short answer in one phrase or single word:" and for TextVQA we used "Question: <question> Short answer in a very single phrase:". For all tasks we used greedy decoding with a maximum generation length of 56 tokens. For each evaluation run we experiment by separating demonstrations using <|endofchunk|> or double newline delimiters. We empirically find that MINT trained models perform better when using double newline delimiters while OBELICS trained model performs better using <|endofchunk|> separators. We do not report the best score for each evaluation trial by ablating over separators but rather choose the overall best prompt for a model based on aggregate scores and report results using that prompt for all evaluations.
@@ A.3.1 Motivation
    The dataset is created by a team from the University of Washington, Salesforce Research, Stanford University, University of Texas at Austin, and University of California Berkeley.
@@ A.3.2 Composition
    In total there are 1057.2 million (M) documents (1029.4M HTML documents, 26.8M PDF documents, and 0.87M ArXiv documents).
@@
    MINT-1T’s HTML documents are filtered from CommonCrawl WARC dumps from 2017 to 2024 based on text quality, the presence of duplicated and undesirable content, and availability of images for downloading. MINT-1T PDF documents are filtered from CommonCrawl WAT dumps from 2023 to 2024 also based on text quality and the presence of duplicated and undesirable content. MINT-1T ArXiv documents is a subset of all ArXiv documents.
@@
    MINT-1T is not self-contained and relies on downloading external image urls to collect the full dataset. There are no guarantees these urls will remain available over time as link rot is a common problem for large scale image datasets. Moreover, as this dataset is over 300TB large, it is infeasible for us to host the full datasets (with images included). There are no restrictions regarding downloading images from these external urls.